Certain fish, such as the Nile fish (Gnathonemus), concentrate charges in their head and tail, thereby producing an electric field in the water around them. This field creates a potential difference of a few volts between the head and tail, which in turn causes current to flow in the conducting seawater. As the fish swims, it passes near objects that have resistivities different from that of seawater, which in turn causes the current to vary. Cells in the skin of the fish are sensitive to this current and can detect changes in it. The changes in the current allow the fish to navigate.Since the electric field is weak far from the fish, we shall consider only the field running directly from the head to the tail. We can model the seawater through which that field passes as a conducting tube of area and having a potential difference across its ends. These fish navigate by responding to changes in the current in seawater. This current is due to a potential difference of around 3.0 V generated by the fish and is about 14 mA within a centimeter or so from the fish. Receptor cells in the fish are sensitive to the current. Since the current is at some distance from the fish, their sensitivity suggests that these cells might be responding to the magnetic field created by the current. To get some estimate of how sensitive the cells are, we can model the current as that of a long, straight wire with the receptor cells 2.00 cm away

What is the strength of the magnetic field at the receptor cells?

Express your answer using two significant figures.

B= ______μT

.Please give a numerical answer!!!!
Answer #1

B = u0I/2pi*r = (4pi*10-7 * 14 * 10-3)/(2pi * 2 * 10-2)

B = 0.14 * 10-6 T = 0.14 uT